log3(x+2) = 1+ (log3x)/2


1
Expert's answer
2021-08-02T13:44:05-0400

"2log_3(x+2)=2+log_3x"

"log_3(x+2)^2=log_39+log_3x"

"log_2(x+2)^2=log_39x"

"(x+2)^2=9x"

"x^2-5x+4=0"

"x_1=1"

"x_2=4"